Break-Even Sales and Sales to Realize Income from Operations

For the current year ended October 31, Friedman Company expects fixed costs of $550,000, a unit variable cost of $52, and a unit selling price of $77.

a. Compute the anticipated break-even sales (units).
units

b. Compute the sales (units) required to realize income from operations of $127,500.
units
Answer #1

a) Break even unit = Fixed cost/Contribution margin per unit = 550000/(77-52) = 22000 Units

b) Required unit = (Fixed cost+desired profit)/Contribution margin per unit = (550000+127500)/25 = 27100 Units